OpenStudy (anonymous):

Suppose that the functions r and s are defined for all real numbers x as follows. R(x)=2x-4 S(x)=x-1 Write the expressions for (r*s)(x)= (r-s)(x)= (r+s)(2)=

OpenStudy (trojanpoem):

(r*s)(x) = R(x) * S(x) (r-s)(x) = R(x) - S(x) (r+s)(x) = R(x) + S(x) Can you continue ?

OpenStudy (anonymous):

2x-4*x-1 2x-4-x-1 2x-4+x-1

OpenStudy (trojanpoem):

Awesome , now do your calculations , use parentheses. (2x-4) (x-1) = 2x^2 -2x -4x + 4 = 2x^2 -6x + 4 (2x-4) - (x-1) = 2x-4 -x +1 = x -3 (2x-4) + (x-1) = 3x - 5

OpenStudy (anonymous):

tyty